If five arithimetic mean are inserted between -9 and 9.what is the third mean?
Answer by amarjeeth123(569)

Let the means be a1,a2,a3,a4,a5.
Then we have -9,a1,a2,a3,a4,a5,9 are in A.P.
-9+6d=9
6d=18
d=3
The third mean is a3=a+3d=-9+3(3)=0
The third mean is zero.